3. The equilibrium constant K, for the formation of COBra, is given below. Kp = 4.7 x 107 at 150°C Answer(b) Calculate Kp for the following reaction at 150°C. (Refer to the reaction and Kp = 4.7 x 107 from the previous page.) (1 p
Answer #1

Part a).

Kp = Kc*(RT)\Deltan

i.e. 4.7*107 = Kc * {0.08206*(150+273.15)}1-2

Therefore, Kc = 2.42*1011